OpenStudy (anonymous):

the width of a soccer field should be 60% of its length. write and simplify an expression for the perimeter of a soccer field with a length of x feet

OpenStudy (whpalmer4):

let W be the width and L be the length. We know that W = 60% of L = 60/100*L = 0.6 L We also know that the perimeter P = 2 * L + 2 * W. Combine the two equations to get an expression for P in terms of L.

OpenStudy (anonymous):

check length is x thus width =60% of x=0.06x now perimeter is given by 2*(length+width) hope getting expression will be easy ...